OpenStudy (anonymous):

HELP....the length of a rectangle is twice the width. the area is 18yd^2. find the length and the width. L= W=

OpenStudy (anonymous):

L=2W then S=L*W=2W*W=2W^2=18=>W^2=9=>W=3 and L=6